### v2

# Leveraging Diverse Semantic-based Audio Pretrained Models for Singing Voice Conversion

Abstract
Singing Voice Conversion (SVC) is a technique that enables any singer to perform any song. To achieve this, it is essential to obtain speaker-agnostic representations from the source audio, which poses a significant challenge. A common solution involves utilizing a semantic-based audio pretrained model as a feature extractor. However, the degree to which the extracted features can meet the SVC requirements remains an open question. This includes their capability to accurately model melody and lyrics, the speaker-independency of their underlying acoustic information, and their robustness for in-the-wild acoustic environments. In this study, we investigate the knowledge within classical semantic-based pretrained models in much detail. We discover that the knowledge of different models is diverse and can be complementary for SVC. To jointly utilize the diverse pretrained models with mismatched time resolutions, we propose an efficient ReTrans strategy to address the feature fusion problem. Based on the above, we design a Singing Voice Conversion framework based on Diverse Semantic-based Feature Fusion (DSFF-SVC). Experimental results demonstrate that DSFF-SVC can be generalized and improve various existing SVC models, particularly in challenging real-world conversion tasks.
1 Introduction
Singing Voice Conversion (SVC) aims to transform the voice of a singing signal into the voice of a target singer while preserving the lyrics and melody Huang et al. (2023). This allows any singer to perform any song. It has a wide range of applications, such as music entertainment, singing voice enhancement, vocal education, and artistic creation.
In recent years, conducting singing voice conversion with non-parallel data  Nachmani and Wolf (2019); Chen et al. (2019) has attracted more attention. Figure 1 displays the classic pipeline for it. To empower the reference speaker to sing the source audio, the main idea is to extract the speaker-specific representations from the reference, extract the speaker-agnostic representations from the source, and then synthesize the converted audio using a decoder. Usually, the speaker-specific representations can be just a one-hot speaker ID Qian et al. (2019) or the features extracted from a pretrained speaker verification model Qian et al. (2019); Nercessian (2020). For speaker-agnostic representations, common solutions involve using the intermediate output [^1] The intermediate output is a dense high-dimensional vector instead of just symbolic token. It is believed that such output contains not only semantic but also acoustic information, which can enhance the quality of synthesized audio. In this paper, we refer to it as semantic-based features. from a semantic-based audio pretrained model. This model’s pretraining task is typically designed to be semantic-related, such as Automatic Speech Recognition (ASR) Chen et al. (2019); Li et al. (2021) or semantic-guided self-supervised learning Huang et al. (2022); Jayashankar et al. (2023).
Figure 1: The role of semantic-based pretrained model in the classic singing voice conversion pipeline.
Requirements of SVC Capability of the Semantic-based Features
To model melody Whether could or not remains unknown
To model lyrics Could. But exactly how much remains unknown
To model auxiliary acoustic information Could. But whether the information is speaker-agnostic or not remains unknown
To be robust for in-the-wild acoustic environment Whether is robust or not remains unknown
Table 1: The extent to which the existing semantic-based features satisfy the requirements of singing voice conversion is still unclear.
For SVC, high-quality speaker-agnostic representations should meet several requirements (Table 1). First, they should be capable of modeling melody and lyrics. Besides, they could also contain some auxiliary acoustic information (such as pronunciation, articulation, and prosody) to improve the naturalness and expressiveness of the synthesized audio. Last but not least, they should be robust for varied acoustic environments of source audios (such as the in-the-wild singing voices separated from background music Takahashi et al. (2022); Zang et al. (2024)). However, for the semantic-based audio pretrained models, our understanding of their underlying knowledge remains limited, despite the considerable resources invested in pretraining Baevski et al. (2020); Hsu et al. (2021); Radford et al. (2023). It is still unclear to what extent the extracted semantic-based features can satisfy the requirements of SVC. For example, except for the semantic signals, there is also acoustic information including prosody in the semantic-based features Wang et al. (2022a); Liu et al. (2023). Is this sufficient, or how much additional information is required to model the melody? Furthermore, is this information speaker-agnostic, or could it cause the source’s timbre leakage Sha et al. (2024) into the converted audio? Moreover, are these semantic-based pretrained models robust to diverse acoustic environments? How effective are the semantic-based features when faced with in-the-wild audio data?
Motivated by this, by exploring the gap when using only semantic-based features to conduct the SVC task, we investigate the underlying knowledge within the three classic semantic-based pretrained models respectively – WeNet Yao et al. (2021), Whisper Radford et al. (2023), and ContentVec Qian et al. (2022). Furthermore, we suppose that the capabilities of semantic-based features mainly depend on the pretraining tasks and data, and different pretraining ways will yield different underlying knowledge (Table 2). Based on this, we propose a concise and effective solution: by utilizing diverse semantic-based models that are pretrained distinctly, the capabilities of these jointly used semantic-based features will be enhanced for SVC. However, it is challenging to fuse the multiple features of diverse audio pretrained models. The reason is that the time resolutions of different models are usually mismatched, since their acoustic parameters like sampling rate and hop size can be different. To address it, we propose a signal processing based resolution transformation (ReTrans) strategy. Given the features of any audio pretrained models, we simply need to apply an upsampling followed by a downsampling to transform these features to our desired time resolution. The experiments verify that this method could achieve the comparable performance to cross attention Vaswani et al. (2017); Shen et al. (2024) but with a lower latency and no additional training cost (Section 4.4).
The contributions of this paper are summarized as follows:
  • We investigate the extent to which the classic semantic-based pretrained models can meet the requirements of SVC. Furthermore, we validate that multiple semantic-based models, pretrained in diverse ways, can complement each other comprehensively under the SVC task.
  • We propose an efficient strategy, ReTrans, to align the time resolutions of features from different audio pretrained models. It can be widely adopted in audio features fusion and multiple pretrained models’ joint usage.
  • Based on the above, we design a Singing Voice Conversion framework based on Diverse Semantic-based Features Fusion (DSFF-SVC). Both objective and subjective evaluation results verify the effectiveness, generalization, and robustness of our proposed framework.
Audio samples are available at our demo site [^2] https://diversesemanticsvc.github.io/ and our code is open-sourced in Amphion [^3] https://github.com/open-mmlab/Amphion Zhang et al. (2024).
2 Related Work
The early singing voice conversion researches aim to design parametric statistical models such as HMM Türk et al. (2009) or GMM Kobayashi et al. (2014, 2015) to learn the spectral features mapping of the parallel data. Since the parallel singing voice corpus is hard to collect on a large scale, the non-parallel SVC Nachmani and Wolf (2019); Chen et al. (2019), or recognition-synthesis SVC Huang et al. (2022), is popular in recent years. The classic pipeline of the non-parallel SVC is displayed in Figure 1. To decouple speaker-agnostic and speaker-specific representations, some pioneering works leverage the information bottleneck Qian et al. (2019) and the adversarial learning Nachmani and Wolf (2019); Deng et al. (2020) to design the end-to-end network. To introduce more explicit guidance instead of depending only on end-to-end learning, utilizing pretrained models as a feature extractor has become the trend.
For the speaker-agnostic representations, the most well-known solution is to leverage the phonetic posteriorgrams (PPG) from pretrained ASR models as the semantic-based features Sun et al. (2016); Li et al. (2021). In recent years, except for the supervised ASR models, there are also more and more semantic-based pretrained models under weak-supervised and self-supervised learning. The researchers have explored in conducting the SVC task based on HuBERT Hsu et al. (2021); Huang et al. (2022), wav2vec 2.0 Baevski et al. (2020); Jayashankar et al. (2023), Whisper Radford et al. (2023); Ning et al. (2023) and more.
However, the specific role of the semantic-based pretrained models and the extent to which the extracted features can meet the SVC requirements is still an open question. For example, some researchers point out that there is also rich acoustic information like prosody in these features Liu et al. (2023), which is beneficial to improve the naturalness and the speaker similarity Wang et al. (2022a). On the contrary, others believe that this acoustic information could be speaker-specific, which will cause the source’s timbre leakage Sha et al. (2024). Motivated by this, we aim to investigate the specific knowledge within the semantic-based features by exploring to use them alone to conduct the SVC task. Besides, most existing works utilize a single source of semantic-based features. This study also researches whether different pretrained models can be complementary to each other for SVC.
3 Methodology
School Representative Pretrained Task Training Objective Pretrained Data Architecture
Supervised WeNet Yao et al. (2021) ASR, supervised by linguistic labels CTC, Next Token Prediction (character-level) Text-only transcription from Gigaspeech/Wenet- speech (10k hours English /Chinese speech) Encoder-decoder Conformer Gulati et al. (2020) or Transformer Vaswani et al. (2017)
Weak-Supervised Whisper Radford et al. (2023) Multitask including multilingual ASR, speech translation, spoken language identi- fication, etc., large-scale weak supervision Next Token Prediction (byte-level Sennrich et al. (2016)) 680k hours multilingual data, including both text- only and time-aligned transcription Encoder-decoder Transformer Vaswani et al. (2017)
Self-Supervised ContentVec Qian et al. (2022) Self-supervised learning which conditions on disentangling speakers Masked Token Prediction (frame-level) Librispeech (1k hours English), using only speech but no any transcription Encoder-only Transformer Vaswani et al. (2017)
Table 2: A systematic analysis for the three schools of the existing semantic-based pretrained models.
In this section, through an analysis of classic semantic-based pretrained models, we first illustrate the potential knowledge in semantic-based features and explain why the joint usage of multiple pretrained models could work for SVC (Section 3.1). Furthermore, we describe how the proposed ReTrans strategy could tackle the features fusion problem for multiple pretrained models with mismatched time resolutions (Section 3.2). Based on the above, we propose the Diverse Semantic-based Features Fusion SVC framework (Figure 2), which is capable of incorporating various models as a base.
3.1 Preliminary Analysis for Semantic-based Pretrained Models
To make a model capture semantic signals of audio, we need to push its latent representations to align with semantic representations (such as characters or phonemes). In other words, the extent to which the latent representations contain semantic information depends on the level of semantic supervision with which the audio model is pretrained. Furthermore, the knowledge of these latent representations that goes beyond semantic information is also determined by the pretraining ways. Based on the above, we categorizes the existing semantic-based pretrained models into three schools (Table 2):
  • Supervised models: The models are pretrained under the ASR task, whose representative is WeNet Yao et al. (2021). For such models, the intermediate layers closer to the input (audio) contain more acoustic information, while layers closer to the output (character) contain more semantic information.
  • Weak-supervised models: The models are pretrained under both ASR task and other auxiliary tasks, whose representative is Whisper Radford et al. (2023). Compared to supervised ones, such models are likely to contain additional knowledge related to the auxiliary tasks.
  • Self-supervised models: The models conduct self-supervised learning under semantic (or speaker-agnostic) guidance, whose representative is ContentVec Qian et al. (2022). The knowledge of these models is highly related to the construction of pseudo labels during the initial audio tokenization, since these labels usually play a role of “teacher” Hsu et al. (2021). Compared to the other two, the self-supervised ones contain more macro and general knowledge that is beneficial for modeling context.
In summary, the underlying knowledge of the three is likely to be different and diverse. In this study, we respectively select a representative for them – WeNet Yao et al. (2021), Whisper Radford et al. (2023), and ContentVec Qian et al. (2022), to explore the gap when adopting them into the SVC task. We will also investigate whether the three can be complementary to each other.
Figure 2: The proposed Singing Voice Conversion framework based on Diverse Semantic-based Features Fusion (DSFF-SVC). It is capable of incorporating most existing models (i.e., acoustic model and waveform decoder) as a base.
3.2 Resolution Transformation for Features of Audio Pretrained Models
For the joint usage of multiple audio pretrained models, there is a technical challenge that the time resolutions of different models are usually mismatched. Specifically, given an utterance, the frame lengths of features extracted by different pretrained models can vary. To align the varied frame lengths, cross attention Vaswani et al. (2017); Shen et al. (2024) can be a classic solution. However, cross attention will introduce a large of parameters and training costs (Section 4.4). To address it, we propose a non-learning and signal processing based Resolution Transformation strategy (ReTrans), which is detailed in Algorithm 1.
Algorithm 1 ReTrans for features of audio pretrained models
Input: $\mathbf{f}$ – the features to be transformed whose frame rate is $r_{1}$ Hz.
Parameter: $r_{2}$ – The desired frame rate to transform
Output: $\mathbf{f}^{{}^{\prime}}$ – the transformed features whose frame rate is $r_{2}$ Hz
1:$c\leftarrow\mathbf{gcd}~{}(r_{1},r_{2})$triangleright The greatest common divisor
2:$\mathbf{f}^{{}^{\prime}}\leftarrow\mathbf{upsample}~{}(\mathbf{f},r_{2}/c)$triangleright Upsampling
3:$\mathbf{f}^{{}^{\prime}}\leftarrow\mathbf{downsample}~{}(\mathbf{f}^{{}^{\prime}},r_{1}/c)$triangleright Downsampling
4:return $\mathbf{f}^{{}^{\prime}}$
According to Algorithm 1, assuming the duration of input audio is $d$ seconds, given the pretrained model whose frame rate is $r_{1}$ Hz, the frame lengths of the extracted features will be $d\times r_{1}$. After an upsampling and subsequent downsampling, the frame lengths of the transformed features will be $d\times r_{2}$. Here we calculate the greatest common divisor first to improve the efficiency of the transformation.
It is worth noting that our proposed ReTrans algorithm can be widely applied to various audio pretrained models. For instance, it can be utilized for the features fusion of multiple pretrained models (like Section 3.3). Another typical scenario is the joint usage of multiple pretrained models, such as jointly using a semantic-based pretrained model and a pretrained vocoder with different time resolutions.
3.3 Singing Voice Conversion Framework based on Diverse Semantic-based Features Fusion
The proposed Singing Voice Conversion framework based on Diverse Semantic-based Features Fusion (DSFF-SVC) is displayed in Figure 2. We leverage the fused multiple features sourcing from diverse semantic-based pretrained models as the speaker-agnostic representations. We adopt a trainable embedding layer to model the speaker features from a one-hot speaker ID. Besides, we can also use the auxiliary prosody features like fundamental frequency (F0) and energy to improve the melody modeling and the expressiveness of synthesized audios.
For the semantic-based pretrained models, the WeNet, Whisper, and ContentVec are all Transformer-based Vaswani et al. (2017) architecture. Here we utilize their encoder’s output as the semantic-based features. Formally, given the utterance $u$ and the acoustic model $\mathcal{M}$, the extracted content features $\mathbf{c}_{\mathcal{M}}=\mathcal{M}(u)\in\mathbb{R}^{T_{\mathcal{M}}\times D_{\mathcal{M}}}$, where $\mathcal{M}$ can be either WeNet, Whisper, or ContentVec, $T_{\mathcal{M}}$ is the frame length, and $D_{\mathcal{M}}$ is the latent representation dimension of the model $\mathcal{M}$. Using the proposed ReTrans, we transform their frame lengths as the same (i.e., $\mathbf{\hat{c}}_{\mathcal{M}}\in\mathbb{R}^{T\times D_{\mathcal{M}}}$), and adopt adding fusion to merge them:
$\mathbf{\hat{c}}=\mathbf{Linear}(\mathbf{\hat{c}}_{\mathcal{M}_{1}})\oplus\mathbf{Linear}(\mathbf{\hat{c}}_{\mathcal{M}_{2}})\oplus\mathbf{Linear}(\mathbf{\hat{c}}_{\mathcal{M}_{3}})$ (1)
where $\mathbf{\hat{c}}\in\mathbb{R}^{T\times D}$, $\mathbf{Linear}$ means the linear layer, $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, and $\mathcal{M}_{3}$ means WeNet, Whisper, and ContentVec, and $\oplus$ means the element-wise adding operation.
For the auxiliary prosody features, we follow Liu et al. (2021) to obtain the quantized F0 and energy features. We adopt the trainable embedding layers to get the F0 embeddings $\mathbf{f}\in\mathbb{R}^{T\times D}$ and energy embeddings $\mathbf{e}\in\mathbb{R}^{T\times D}$. For the speaker-specific representations, we adopt a look-up table and use a trainable embedding layer to learn the speaker embeddings $\mathbf{s}\in\mathbb{R}^{D}$. Finally, we adopt the adding fusion to merge all the features:
$\mathbf{c}=\mathbf{CondEnc}(\mathbf{\hat{c}}\oplus\mathbf{f}\oplus\mathbf{e}\oplus\mathbf{\hat{s}})$ (2)
where $\mathbf{CondEnc}$ means a condition encoder that interacts with all the conditions of SVC. It is set as a simple linear layer in our experiments. $\mathbf{\hat{s}}\in\mathbb{R}^{T\times D}$ represents the frame-level speaker feature, being obtained by repeating $\mathbf{s}$ $T$ times.
During training, the DSFF-SVC conducts the reconstruction learning on the training corpus, and the speaker ID is just the speaker identity of every training sample. During conversion, given any source audio, we extract its semantic-based and energy features and stay them unchanged. To convert the speaker identity, we inject a reference speaker ID (which is seen in the training) to obtain the speaker-specific representations. For F0 features, we conduct the musical key transposition [^4] https://en.wikipedia.org/wiki/Transposition_(music) to make the reference speaker sing the source song in his vocal range. Specifically, following Liu et al. (2021); Huang et al. (2023), we shift the source F0 features by multiplying a factor, which is computed as the ratio between the F0 medians among the reference speaker’s training corpus and the source audio.
Notably, the DSFF-SVC framework can support any architectures of acoustic models and waveform decoders (vocoders). During our experiments, we will investigate the generalization of DSFF-SVC for various base models including transformer-based, diffusion-based, and end-to-end models (Section 4.1).
4 Experiments
We conduct experiments to answer the following evaluation questions:
  • EQ1: How much the existing semantic-based pretrained models can meet the requirements of SVC? Could the multiple features from diverse semantic-based pretrained models be complementary?
  • EQ2: How effective and generalized is the proposed singing voice conversion framework based on diverse semantic-based features fusion?
  • EQ3: How effective is the proposed resolution transformation strategy for fusing multiple features of mismatched time resolutions?
4.1 Experimental Setup
4.1.1 Evaluation Tasks
We adopt two conversion settings to evaluate: (1) Recording Studio Setting: following the most existing works, we utilize the high-quality singing corpus that is recorded in studio as the experimental data. The vocals are clean with virtually no noise or environmental interference. Specifically, we use Opencpop Wang et al. (2022b) as the target singer, whose training corpus is 5.2 hours of studio recorded singing voices. For source audios, we use M4Singer Zhang et al. (2022) and randomly 25 utterances for each timbre type respectively (including Soprano, Alto, Tenor, and Bass). (2) In-the-Wild Setting: in the real-world SVC application, usually the singing voices are separated from the background music, which will remain some artifacts or reverb in the vocals Takahashi et al. (2022); Zang et al. (2024). We consider this as a more challenging conversion task to examine the robustness of the SVC systems. Specifically, we adopt a private corpus which contains 6.4 hours of singing voices of 15 professional singers (6 English singers and 9 Chinese singers). The vocals are separated by Ultimate Vocal Remover (UVR) [^5] https://github.com/Anjok07/ultimatevocalremovergui. We adopt four singers as the targets (an English male, an English female, a Chinese male, and a Chinese female). For source audios, we randomly sample 100 utterances which cover multiple musical genres including Pop, Rock, Folk and Soul.
4.1.2 Evaluation Metrics
For objective evaluation, following Huang et al. (2023), we adopt Mel-cepstral distortion (MCD) Kubichek (1993), F0 Pearson correlation coefficient (F0CORR), F0 Root Mean Square Error (F0RMSE), Character Error Rate (CER) which is obtained with the recognition results of whisper-large ASR model Radford et al. (2023). Besides, following Jiang et al. (2023), we use the WavLM model Chen et al. (2022) finetuned for speaker verification [^6] https://huggingface.co/microsoft/wavlm-base-plus-sv to compute the cosine speaker similarity score (SIM). For subjectivce evalution, we invite 12 volunteers who are experienced in the audio generation areas to conduct the Mean Opinion Score (MOS) evaluation in terms of naturalness and similarity. The naturalness score ranks from 1 (“Bad”) to 5 (“Excellent”), and the similarity score ranks from 1 (“Different speaker, sure”) to 4 (“Same speaker, sure”).
4.1.3 Base Models
We select three base models to verify the generalization ability of DSFF-SVC framework:
  • TransformerSVC: It adopts a vanilla encoder-only transformer model Vaswani et al. (2017) as the acoustic model. Its output is the mel-spectrogram.
  • VitsSVC: It is a VITS-based Kim et al. (2021) model which is similar to the SoftVC-VITS [^7] https://github.com/svc-develop-team/so-vits-svc. It is an end-to-end framework and can directly produce waveform.
  • DiffWaveNetSVC: It adopts a diffusion-based acoustic model and could generate mel-spectrogram, which is proposed by Liu et al. (2021). The internal encoder of the diffusion framework is based on Bidirectional Non-Causal Dilated CNN Kong et al. (2021), which is similar to WaveNet van den Oord et al. (2016).
4.1.4 Implementation Details
For WeNet, we use the official models pretrained by 10k hours Wenetspeech to extract semantic-based features. For Whisper, we use the multilingual MEDIUM model. For ContentVec, we use the official 500-CLASS model preatrained by 1k hours Librispeech. The latent space dimension $D$ is set as 384. We adopt Parselmouth [^8] https://parselmouth.readthedocs.io/en/stable/index.html Jadoul et al. (2018) to extract F0 and compute L2-norm of the amplitude of each short-time Fourier transform frame as energy features Ren et al. (2021). Following Liu et al. (2021), we set their numbers of bins for quantization as 256. For the three base models, we adopt the implementations of Amphion [^9] https://github.com/open-mmlab/Amphion/tree/main/egs/svc Zhang et al. (2024). For DiffWaveNetSVC and TransformerSVC, we use the pretrained Amphion Singing BigVGAN [^10] https://huggingface.co/amphion/BigVGAN_singing_bigdata as the vocoder to decode waveform from mel-spectrogram.
4.2 Performance of Different Semantic-based Features (EQ1)
In this section, we aim to explore the gap when using semantic-based features alone to conduct the SVC task. We select DiffWaveNetSVC as the base model and conduct the task under the Recording Studio Setting. The experimental results are illustrated in Table 3.
Semantic-based Features MCD ($\downarrow$) F0CORR ($\uparrow$) F0RMSE ($\downarrow$) CER ($\downarrow$) SIM ($\uparrow$)
Ground Truth 0.000 1.000 0.0 12.9% 1.000
WeNet 10.324 0.203 423.4 38.2% 0.912
Whisper 8.229 0.524 297.3 18.9% 0.914
ContentVec 8.972 0.491 361.0 22.1% 0.918
WeNet + Whisper 8.345 0.540 284.2 16.8% 0.911
WeNet + ContentVec 8.870 0.525 329.5 19.9% 0.912
Whisper + ContentVec 8.201 0.548 279.6 16.9% 0.912
WeNet + Whisper + ContentVec 8.249 0.572 278.5 16.1% 0.913
Table 3: Objective evaluation results of different semantic-based features and their integration. It can be observed that from WeNet to Wenet + Whisper, and then to Wenet + Whisper + ContentVec, the results of most metrics are promoted stage by stage.
(a) WeNet
(b) WeNet + Whisper
(c) WeNet + Whisper + ContentVec
Figure 3: The complementary role of diverse semantic-based features in melody modeling.
It can be observed that on the one hand, the underlying knowledge of different semantic-based models can be distinct and diverse: (1) To model melody (F0CORR and F0RMSE), we find that Whisper and ContentVec behave better than WeNet. This is mainly because that the auxiliary tasks of the weak-supervised Whisper and the self-supervised ContentVec will be beneficial for modeling more prosody-related signals. (2) To model lyrics (CER), Whisper is the best compared to the other two. It is supposed that the multilingual ASR tasks and the large-scale pretraining data enable Whisper to model more valid and robust semantic information. Besides, ContentVec will provide better semantic signals than WeNet, although the two are both pretrained on only speech. It reveals that the self-supervised pretrained models could be more robust than the classic supervised ASR model. (3) To measure the speaker-independency (SIM), the speaker similarity results of the three are comparable and all over 0.9, which is hard to rank from human perception (see our demo page). It means that when using only semantic-based features for SVC, the three can all be considered speaker-agnostic.
On the other hand, the diverse semantic-based features can be complementary in most cases. For example, from WeNet to Wenet + Whisper, and then to Wenet + Whisper + ContentVec, the results of most metrics are promoted stage by stage. We display a case study of melody modeling in Figure 3. It illustrates that after integrating diverse semantic-based features, the trajectories of the melody between converted audios and ground truth are closer. However, we can also find that using only semantic-based features is hard to model melody adequately (the highest F0CORR in Table 3 is 0.572), appearing the “out of tune” for human hearing. Therefore, introducing explicit melody modeling (such as F0 features) for SVC remains necessary in the present technology context.
4.3 Performance of the DSFF-SVC framework (EQ2)
To verify the generalization and robustness of the proposed DSFF-SVC framework, we conduct the both recording studio and in-the-wild setting tasks for the three base models. Except for the semantic-based features, here we also use the auxiliary prosody features (including F0 and energy), to improve the melody modeling and the expressiveness of the system. The objective and subjective evaluation results can be seen in Table 4. It reveals that: (1) Conducting SVC under in-the-wild setting is more difficult than recording studio setting. Facing to the more complex acoustic environment, only using the supervised WeNet produces a very poor performance. Integrating with the weak-supervised Whisper and the self-supervised ContentVec are usually beneficial, especially for Whisper. (2) For the generalization ability, we can observe that for all the three base models, the proposed idea of diversion semantic-based features fusion is effective in most cases under both recording studio and in-the-wild settings. (3) For the robustness, particularly, under the more challenging in-the-wild conversion setting, the improvement of integrating diverse semantic-based features is more obvious, especially for the objective CER and the subjective metrics.
Base Model Semantic-based Features Recording Studio Setting In-the-Wild Setting
F0CORR ($\uparrow$) F0RMSE ($\downarrow$) CER ($\downarrow$) SIM ($\uparrow$) F0CORR ($\uparrow$) F0RMSE ($\downarrow$) CER ($\downarrow$) SIM ($\uparrow$)
TransformerSVC WeNet 0.849 149.3 15.6% 0.878 0.871 210.0 40.0% 0.865
WeNet + Whisper 0.924 77.2 14.9% 0.881 0.848 183.8 18.7% 0.867
WeNet + Whisper + ContentVec 0.931 75.5 16.2% 0.883 0.857 186.7 23.3% 0.868
VitsSVC WeNet 0.937 175.3 19.1% 0.890 0.919 91.3 57.7% 0.869
WeNet + Whisper 0.945 144.4 17.8% 0.890 0.920 86.9 35.2% 0.869
WeNet + Whisper + ContentVec 0.946 112.9 17.7% 0.886 0.921 79.5 32.3% 0.870
DiffWaveNetSVC WeNet 0.936 55.5 15.8% 0.875 0.901 87.8 60.8% 0.855
WeNet + Whisper 0.943 49.5 15.2% 0.884 0.921 73.6 21.1% 0.865
WeNet + Whisper + ContentVec 0.940 55.2 15.7% 0.884 0.919 79.9 23.3% 0.867
(a) Objective evaluation results.
Semantic-based Features Recording Studio Setting In-the-Wild Setting
Naturalness ($\uparrow$) Similarity ($\uparrow$) Naturalness ($\uparrow$) Similarity ($\uparrow$)
WeNet 2.72 ±$\pm 0.22$ 2.64 ±$\pm 0.21$ 2.85 ±$\pm 0.21$ 2.34 ±$\pm 0.20$
WeNet + Whisper 4.02 ±$\pm 0.18$ 3.13 ±$\pm 0.17$ 3.70 ±$\pm 0.18$ 2.86 ±$\pm 0.23$
WeNet + Whisper + ContentVec 4.14 ±$\pm 0.19$ 3.25 ±$\pm 0.18$ 3.71 ±$\pm 0.18$ 2.82 ±$\pm 0.23$
(b) MOS evaluation results (with 95% confidence interval) of DiffWaveNetSVC. The full scores of Naturalness and Similarity are 5 and 4.
Table 4: Results of the proposed DSFF-SVC framework.
Fusion Strategy Computational Cost Conversion Quality
RTX ($\uparrow$) RTF ($\downarrow$) F0CORR ($\uparrow$) F0RMSE ($\downarrow$) CER ($\downarrow$) SIM ($\uparrow$)
Cross attention 191.9 2.57 0.896 88.6 16.4% 0.871
ReTrans 415.0 0.86 0.936 55.5 15.8% 0.875
Table 5: The comparison between the proposed ReTrans and cross attention for features fusion.
Compared with Table 3 and 4, there is also many interesting observations about introducing the auxiliary prosody features (F0 and energy): (1) After injecting such prosody signals into SVC models, the intelligibility (CER) has been improved a lot. For example, using WeNet features on DiffWaveNetSVC, the CER is decreased from 38.2% to 15.8%. In our opinions, this is because that the prosody in singing voice are expressive and characteristic, which is hard to be modeled without the explicit assistant of prosody features like F0. Therefore, using only semantic-based features, some tone- and intonation-related signals will not be learned well, leading to the worse intelligibility. (2) Except for the intelligibility, it is also notable about the change of conversion speaker similarity (SIM). After introduce the explicit assistant of F0 features, the melody-related metrics have been improved a lot and the converted singing voices are not out of tune any more. In the meantime, however, the speaker similarity is also harmed. We suppose that such auxiliary prosody features own some speaker-specific signals, such as the personalized vibrato patterns within F0 features. Therefore, just copying the original prosody features from the source audio could not be the best choice for SVC, although it is a common way in the most SVC works.
4.4 Performance of the Resolution Transformation based Features Fusion (EQ3)
To verify the necessity and efficiency of the proposed ReTrans strategy for feature alignment of mismatched time resolutions, we select the classic cross attention Vaswani et al. (2017) as a baseline. Specifically, under DSFF-SVC framework, we utilize only WeNet to extract semantic-based features and try different strategies to align its resolution to that of F0 and energy features. We conduct the experiments on DiffWaveNetSVC and follow Shen et al. (2024) to adopt the cross attention for features fusion within the diffusion model.
Except for the quality evaluation for the conversion results, we also measure the computational cost of the both. Assume the duration of the whole training corpus is $d_{train}$, the duration of the evaluation samples to be converted is $d_{infer}$, the time to train an epoch is $t_{train}$, and the time to infer the evaluation samples is $t_{infer}$. We define RTX as $d_{train}/t_{train}$, which can measure the training efficiency. We define RTF as $t_{infer}/d_{infer}$, which can measure the inference latency. The computational platform is a single NVIDIA Tesla V100.
The comparison results between ReTrans and cross attention are displayed in Table 5. We can see that when fusing the features of mismatched time resolutions for the SVC task, our proposed ReTrans is slightly better than cross attention in terms of conversion quality. However, as a non-learning method, ReTrans does not introduce any additional parameters and greatly accelerates both training efficiency (RTX) and inference speed (RTF) much.
5 Conclusion and Future Work
This paper investigates the underlying knowledge within the three representatives from supervised, weak-supervised, and self-supervised semantic-based pretrained models for singing voice conversion. We discover that the capabilities including modeling melody and lyrics of different semantic-based features are diverse and can be complementary. Based on these findings, we propose to utilize diverse semantic-based pretrained models to help each other for SVC. Besides, for joint usage of multiple models under mismatched time resolutions, we design the ReTrans strategy which owns high training efficiency and low inference latency. The experimental results verify the effectiveness of our proposed SVC framework based on Diverse Semantic-based Features Fusion, especially for the in-the-wild data of real-world applications.
In addition, this paper raises other questions worth exploring. Is it possible to pretrain one (or several) model(s) that can capture both melody and lyrics signals while being speaker-agnostic at the same time? How should we select the pretraining method and construct the pretraining corpus to improve the robustness? We will leave these to our future research.