### v3

# Leveraging Diverse Semantic-based Audio Pretrained Models for Singing Voice Conversion

Abstract
Singing Voice Conversion (SVC) is a technique that enables any singer to perform any song. To achieve this, it is essential to obtain speaker-agnostic representations from the source audio, which poses a significant challenge. A common solution involves utilizing a semantic-based audio pretrained model as a feature extractor. However, the degree to which the extracted features can meet the SVC requirements remains an open question. This includes their capability to accurately model melody and lyrics, the speaker-independency of their underlying acoustic information, and their robustness for in-the-wild acoustic environments. In this study, we investigate the knowledge within classical semantic-based pretrained models in much detail. We discover that the knowledge of different models is diverse and can be complementary for SVC. Based on the above, we design a Singing Voice Conversion framework based on Diverse Semantic-based Feature Fusion (DSFF-SVC). Experimental results demonstrate that DSFF-SVC can be generalized and improve various existing SVC models, particularly in challenging real-world conversion tasks. Our demo website is available at https://diversesemanticsvc.github.io/.
Index Terms—  Singing Voice Conversion, Semantic Features, Features Fusion, Robustness
1 Introduction
Singing Voice Conversion (SVC) aims to transform a singing signal into the voice of a target singer while maintaining the original lyrics and melody [1]. This allows any singer to perform any song. It has a wide range of applications, such as music entertainment, singing voice enhancement, vocal education, and artistic creation.
In recent years, generative models [2, 3] and conducting singing voice conversion with non-parallel data  [4, 5] has attracted more attention. Figure 1 displays the classic pipeline for it. To empower the reference speaker to sing the source audio, the main idea is to extract the speaker-specific representations from the reference, extract the speaker-agnostic representations from the source, and then synthesize the converted audio using a decoder. Usually, speaker-specific representations can be just a one-hot speaker ID [6] or features extracted from a pretrained speaker verification model [6, 7]. For speaker-agnostic representations, common solutions involve using the intermediate output [^1] The intermediate output is a dense high-dimensional vector instead of just symbolic token. It is believed that such output contains not only semantic but also acoustic information, which can enhance the quality of synthesized audio. In this paper, we refer to it as semantic-based features. from a semantic-based audio pretrained model. The pretraining task of this model is typically designed to be semantic-related, such as Automatic Speech Recognition (ASR) [5, 8] or self-supervised learning guided by semantics [9, 10].
Fig. 1: The role of semantic-based pretrained model in the classic singing voice conversion pipeline.
Requirements of SVC Capability of the Semantic-based Features
To model melody Whether could or not remains unknown
To model lyrics Could. But exactly how much remains unknown
To model auxiliary acoustic information Could. But whether the information is speaker-agnostic or not remains unknown
To be robust for in-the-wild acoustic environment Whether is robust or not remains unknown
Table 1: The extent to which the existing semantic-based features satisfy the requirements of singing voice conversion is still unclear.
For SVC, high-quality speaker-agnostic representations should meet several requirements (Table 1). First, they should be capable of modeling melody and lyrics. Besides, they could also contain some auxiliary acoustic information (such as pronunciation, articulation, and prosody) to improve the naturalness and expressiveness of the synthesized audio. Last, they should be robust for varied acoustic environments of source audios (such as the in-the-wild singing voices separated from background music [11, 12]). However, for the semantic-based audio pretrained models, our understanding of their underlying knowledge remains limited, despite the considerable resources invested in pretraining [13, 14, 15]. It is still unclear to what extent the extracted semantic-based features can satisfy the requirements of SVC. For example, except for the semantic signals, there is also acoustic information including prosody in the semantic-based features [16, 17]. Is this sufficient, or how much additional information is required to model the melody? Furthermore, is this information speaker-agnostic, or could it cause the source’s timbre leakage [18] into the converted audio? Moreover, are these semantic-based pretrained models robust to diverse environments? How effective are these features when faced with in-the-wild audio data?
Motivated by this, by exploring the gap using only semantic-based features to conduct the SVC task, we investigate the underlying knowledge within the three classic semantic-based pretrained models respectively – WeNet [19], Whisper [15], and ContentVec [20]. Furthermore, we suppose that the capabilities of semantic-based features mainly depend on the pretraining tasks and data, and different pretraining ways will yield different underlying knowledge (Table 2). Based on this, we propose a concise and effective solution: by utilizing diverse semantic-based models that are pretrained distinctly, the capabilities of these jointly used semantic-based features will be enhanced for SVC. However, it is challenging to fuse the multiple features of diverse audio pretrained models. The reason is that the time resolutions of different models are usually mismatched, since their acoustic parameters like sampling rate and hop size can be different. To address it, we explored the efficacy and efficiency between a resampling strategy with cross attention [21, 22]. The experiments verify that such signal processing based method could achieve the comparable performance to cross attention but with a lower latency and no additional training cost (Section 4.4). Based on the above, we design a Singing Voice Conversion framework based on Diverse Semantic-based Features Fusion (DSFF-SVC). Both objective and subjective evaluation results verify the effectiveness, generalization, and robustness of our proposed framework.
School Representative Pretrained Task Training Objective Pretrained Data Architecture
Supervised WeNet [19] ASR, supervised by linguistic labels CTC, Next Token Prediction (character-level) Text-only transcription from Gigaspeech/Wenet- speech (10k hours English /Chinese speech) Encoder-decoder Conformer [23] or Transformer [21]
Weak-Supervised Whisper [15] Multitask including multilingual ASR, speech translation, spoken language identi- fication, etc., large-scale weak supervision Next Token Prediction (byte-level [24]) 680k hours multilingual data, including both text- only and time-aligned transcription Encoder-decoder Transformer [21]
Self-Supervised ContentVec [20] Self-supervised learning which conditions on disentangling speakers Masked Token Prediction (frame-level) Librispeech (1k hours English), using only speech but no any transcription Encoder-only Transformer [21]
Table 2: A systematic analysis for the three schools of the existing semantic-based pretrained models.
2 Related Work
The early singing voice conversion researches aim to design parametric statistical models such as HMM [25] or GMM [26, 27] to learn the spectral features mapping of the parallel data. Since the parallel singing voice corpus is hard to collect on a large scale, the non-parallel SVC [4, 5], or recognition-synthesis SVC [9], is popular in recent years. The classic pipeline of the non-parallel SVC is displayed in Figure 1. To decouple speaker-agnostic and speaker-specific representations, some pioneering works leverage the information bottleneck [6] and the adversarial learning [4, 28] to design the end-to-end network. To introduce more explicit guidance instead of depending only on end-to-end learning, utilizing pretrained models as a feature extractor has become the trend.
For the speaker-agnostic representations, the most well-known solution is to leverage the phonetic posteriorgrams (PPG) from pretrained ASR models as the semantic-based features [29, 8]. In recent times, besides the supervised ASR models, an increasing number of semantic-based pretrained models have emerged under weak supervision and self-supervised learning. Researchers have explored in conducting the SVC task based on HuBERT [14, 9], wav2vec 2.0 [13, 10], Whisper [15, 30] and more.
However, the specific role of the semantic-based pretrained models and the extent to which the extracted features can meet the SVC requirements is still an open question. For example, some researchers point out that there is also rich acoustic information like prosody in these features [17], which is beneficial to improve the naturalness and the speaker similarity [16]. On the contrary, others believe that this acoustic information could be speaker-specific, which will cause the source’s timbre leakage [18]. Motivated by this, we aim to investigate the specific knowledge within the semantic-based features by exploring to use them alone to conduct the SVC task. Besides, most existing works utilize a single source of semantic-based features. This study also researches whether different pretrained models can be complementary for SVC.
3 Methodology
Fig. 2: The proposed Singing Voice Conversion framework based on Diverse Semantic-based Features Fusion (DSFF-SVC). It is capable of incorporating most existing models (i.e., acoustic model and waveform decoder) as a base.
In this section, we analyze classic semantic-based pretrained models to demonstrate the potential knowledge embedded in semantic-based features. We also explain why combining multiple pretrained models can be effective for SVC (Section 3.1). Additionally, we discuss the use of resampling and cross-attention strategies for addressing the feature fusion issue of multiple pretrained models with different time resolutions (Section 3.2). Building on this, we introduce the Diverse Semantic-based Features Fusion SVC framework (DSFF-SVC, Figure 2), which can integrate various models as its foundation.
3.1 Analysis for Semantic-based Pretrained Models
In order for a model to capture the semantic signals of audio, we need to push its latent representations to align with semantic representations (such as characters or phonemes). In other words, the extent to which the latent representations contain semantic information depends on the level of semantic supervision with which the audio model is pretrained. Furthermore, the knowledge of these latent representations that goes beyond semantic information is also determined by the pretraining ways. Based on the above, we categorizes the existing semantic-based pretrained models into three schools (Table 2):
  • Supervised models: The models are pretrained under the ASR task, whose representative is WeNet [19]. For such models, the intermediate layers closer to the input (audio) contain more acoustic information, while layers closer to the output (character) contain more semantic information.
  • Weak-supervised models: The models are pretrained under both ASR task and other auxiliary tasks, whose representative is Whisper [15]. Compared to supervised ones, such models are likely to contain additional knowledge related to the auxiliary tasks.
  • Self-supervised models: The models conduct self-supervised learning under semantic (or speaker-agnostic) guidance, whose representative is ContentVec [20]. The knowledge of these models is highly related to the construction of pseudo labels during the initial audio tokenization, since these labels usually play a role of “teacher” [14]. Compared to the other two, the self-supervised ones contain more macro and general knowledge that is beneficial for modeling context.
In summary, the underlying knowledge of the three are likely to be different and diverse. In this study, we respectively select a representative for them – WeNet [19], Whisper [15], and ContentVec [20], to explore the gap when adopting them into the SVC task. We will also investigate whether they can be complementary to each other.
3.2 Features Fusion for Models of Mismatched Resolutions
When combining several pretrained audio models, a technical problem emerges because of the differing time resolutions of the models. In particular, for a specific utterance, the feature frame lengths extracted by different semantic-based pretrained models may vary. A classic signal processing based method for aligning features with different frame rates is to apply resampling. Besides of fusing the different semantic features, another common use case is to integrate a semantic-based model with a pretrained vocoder operating at different time resolutions.
3.3 Singing Voice Conversion Framework based on Diverse Semantic-based Features Fusion
The proposed Singing Voice Conversion framework based on Diverse Semantic-based Features Fusion (DSFF-SVC) is displayed in Figure 2. The fusion of multiple features derived from diverse semantic-based pretrained models serves as the speaker-agnostic representation. A trainable embedding layer is employed to model the speaker features derived from a one-hot speaker ID. Furthermore, the incorporation of auxiliary prosody features, such as fundamental frequency (F0) and energy, facilitates the enhancement of melody modeling and the expressiveness of synthesized audio.
For the semantic-based pretrained models, the WeNet, Whisper, and ContentVec are all Transformer-based [21] architecture. Here we utilize their encoder’s output as the semantic-based features. Formally, given the utterance $u$ and the acoustic model $\mathcal{M}$, the extracted content features $\mathbf{c}_{\mathcal{M}}=\mathcal{M}(u)\in\mathbb{R}^{T_{\mathcal{M}}\times D_{\mathcal{M}}}$, where $\mathcal{M}$ can be either WeNet, Whisper, or ContentVec, $T_{\mathcal{M}}$ is the frame length, and $D_{\mathcal{M}}$ is the latent representation dimension of the model $\mathcal{M}$. Using resampling, we transform their frame lengths as the same (i.e., $\mathbf{\hat{c}}_{\mathcal{M}}\in\mathbb{R}^{T\times D_{\mathcal{M}}}$), and adopt adding fusion to merge them:
$\mathbf{\hat{c}}=\mathbf{Linear}(\mathbf{\hat{c}}_{\mathcal{M}_{1}})\oplus\mathbf{Linear}(\mathbf{\hat{c}}_{\mathcal{M}_{2}})\oplus\mathbf{Linear}(\mathbf{\hat{c}}_{\mathcal{M}_{3}})$ (1)
where $\mathbf{\hat{c}}\in\mathbb{R}^{T\times D}$, $\mathbf{Linear}$ means the linear layer, $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, and $\mathcal{M}_{3}$ means WeNet, Whisper, and ContentVec, and $\oplus$ means the element-wise adding operation.
For the auxiliary prosody features, we follow [31] to obtain the quantized F0 and energy features. We adopt the trainable embedding layers to get the F0 embeddings $\mathbf{f}\in\mathbb{R}^{T\times D}$ and energy embeddings $\mathbf{e}\in\mathbb{R}^{T\times D}$. For the speaker-specific representations, a look-up table with a trainable embedding layer is adopted to learn the speaker embeddings $\mathbf{s}\in\mathbb{R}^{D}$. Finally, we adopt the adding fusion to merge all the features:
Semantic-based Features MCD ($\downarrow$) F0CORR ($\uparrow$) F0RMSE ($\downarrow$) CER ($\downarrow$) SIM ($\uparrow$)
Ground Truth 0.000 1.000 0.0 12.9% 1.000
WeNet 10.324 0.203 423.4 38.2% 0.912
Whisper 8.229 0.524 297.3 18.9% 0.914
ContentVec 8.972 0.491 361.0 22.1% 0.918
WeNet + Whisper 8.345 0.540 284.2 16.8% 0.911
WeNet + ContentVec 8.870 0.525 329.5 19.9% 0.912
Whisper + ContentVec 8.201 0.548 279.6 16.9% 0.912
WeNet + Whisper + ContentVec 8.249 0.572 278.5 16.1% 0.913
Table 3: Objective evaluation results of different semantic-based features and their integration. It can be observed that from WeNet to Wenet + Whisper, and then to Wenet + Whisper + ContentVec, the results of most metrics are promoted stage by stage.
$\mathbf{c}=\mathbf{CondEnc}(\mathbf{\hat{c}}\oplus\mathbf{f}\oplus\mathbf{e}\oplus\mathbf{\hat{s}})$ (2)
where $\mathbf{CondEnc}$ means a condition encoder that interacts with all the conditions of SVC. It is set as a simple linear layer in our experiments. $\mathbf{\hat{s}}\in\mathbb{R}^{T\times D}$ represents the frame-level speaker feature, being obtained by repeating $\mathbf{s}$ $T$ times.
During training, the DSFF-SVC conducts the reconstruction learning on the training corpus, and the speaker ID is just the speaker identity of every training sample. During conversion, given any source audio, we extract its semantic-based and energy features and stay them unchanged. To convert the speaker identity, we inject a reference speaker ID (which is seen in the training) to obtain the speaker-specific representations. For F0 features, we conduct the musical key transposition [^2] https://en.wikipedia.org/wiki/Transposition_(music) to make the reference speaker sing the source song in his vocal range. Specifically, following [31, 1], we shift the source F0 features by multiplying a factor, which is computed as the ratio between the F0 medians among the reference speaker’s training corpus and the source audio.
Notably, the DSFF-SVC framework can support any architectures of acoustic models and waveform decoders (vocoders). During our experiments, we will investigate the generalization of DSFF-SVC for various base models including transformer-based, diffusion-based, and end-to-end models (Section 4.1).
4 Experiments
We conduct experiments to answer the following questions:
  • EQ1: How much the existing semantic-based pretrained models can meet the requirements of SVC? Could the multiple features from diverse semantic-based pretrained models be complementary?
  • EQ2: How effective and generalized is the proposed singing voice conversion framework based on diverse semantic-based features fusion?
  • EQ3: How effective is the resampling strategy compared to deep learning based method such as cross attention [21, 22] for fusing multiple features of mismatched time resolutions?
4.1 Experimental Setup
4.1.1 Evaluation Tasks
We adopt two conversion settings to evaluate: (1) Recording Studio Setting: following the most existing works, we utilize the high-quality singing corpus that is recorded in studio as the experimental data. The vocals are clean with virtually no noise or environmental interference. Specifically, we use Opencpop [32] as the target singer, whose training corpus is 5.2 hours of studio recorded singing voices. For source audios, we use M4Singer [33] and randomly 25 utterances for each timbre type respectively (including Soprano, Alto, Tenor, and Bass). (2) In-the-Wild Setting: in the real-world SVC application, usually the singing voices are separated from the background music, which will remain some artifacts or reverb in the vocals [11, 12]. We consider this as a more challenging conversion task to examine the robustness of the SVC systems. Specifically, we adopt a private corpus which contains 6.4 hours of singing voices of 15 professional singers (6 English singers and 9 Chinese singers). The vocals are separated by Ultimate Vocal Remover (UVR) [^3] https://github.com/Anjok07/ultimatevocalremovergui. We adopt four singers as the targets (an English male, an English female, a Chinese male, and a Chinese female). For source audios, we randomly sample 100 utterances which cover multiple musical genres including Pop, Rock, Folk and Soul.
4.1.2 Evaluation Metrics
For objective evaluation, following [1], we adopt Mel-cepstral distortion (MCD) [34], F0 Pearson correlation coefficient (F0CORR), F0 Root Mean Square Error (F0RMSE), Character Error Rate (CER) which is obtained with the recognition results of whisper-large ASR model [15]. Besides, following [35], we use the WavLM model [36] finetuned for speaker verification [^4] https://huggingface.co/microsoft/wavlm-base-plus-sv to compute the cosine speaker similarity score (SIM). For subjectivce evalution, we invite 12 volunteers who are experienced in the audio generation areas to conduct the Mean Opinion Score (MOS) evaluation in terms of naturalness and similarity. The naturalness score ranks from 1 (“Bad”) to 5 (“Excellent”), and the similarity score ranks from 1 (“Different speaker, sure”) to 4 (“Same speaker, sure”).
4.1.3 Base Models
We select three base models to verify the generalization ability of DSFF-SVC framework:
  • TransformerSVC: It adopts a vanilla encoder-only transformer model[21] as the acoustic model. Its output is the mel-spectrogram.
  • VitsSVC: It is a VITS-based [37] model which is similar to the SoftVC-VITS [^5] https://github.com/svc-develop-team/so-vits-svc. It is an end-to-end framework and can directly produce waveform.
  • DiffWaveNetSVC: It adopts a diffusion-based acoustic model and could generate mel-spectrogram, which is proposed by [31]. The internal encoder of the diffusion framework is based on Bidirectional Non-Causal Dilated CNN[38], which is similar to WaveNet[39].
4.1.4 Implementation Details
For WeNet, we use the official models pretrained by 10k hours Wenetspeech [^6] https://github.com/wenet-e2e/wenet to extract semantic-based features. For Whisper, we use the multilingual MEDIUM model [^7] https://github.com/openai/whisper. For ContentVec, we use the official 500-CLASS model preatrained by 1k hours Librispeech [^8] https://github.com/auspicious3000/contentvec. The latent space dimension $D$ is set as 384. We adopt Parselmouth [40] to extract F0 and compute L2-norm of the amplitude of each short-time Fourier transform frame as energy features [41]. Following [31], we set their numbers of bins for quantization as 256. For the three base models, we adopt the implementations of Amphion [^9] https://github.com/open-mmlab/Amphion/tree/main/egs/svc [42]. For DiffWaveNetSVC and TransformerSVC, we use the pretrained Amphion Singing BigVGAN [^10] https://huggingface.co/amphion/BigVGAN_singing_bigdata as the vocoder to decode waveform from mel-spectrogram.
4.2 Performance of Different Semantic-based Features (EQ1)
In this section, we aim to explore the gap using semantic-based features alone to conduct the SVC task. We select DiffWaveNetSVC as the base model and conduct the task under the Recording Studio Setting. The experimental results are illustrated in Table 3.
(a) WeNet
(b) WeNet + Whisper
(c) WeNet + Whisper + ContentVec
Fig. 3: The complementary role of diverse semantic-based features in melody modeling. More benefits of the joint usage of diverse semantic-based features (including spectrogram reconstruction, lyrics modeling, etc.) can be seen at our demo website.
On the one hand, the underlying knowledge of different semantic-based models can be distinct and diverse: (1) For modeling melody (F0CORR and F0RMSE), Whisper and ContentVec perform better than WeNet. This is mainly because the auxiliary tasks of the weak-supervised Whisper and the self-supervised ContentVec are beneficial for modeling more prosody-related signals. (2) For modeling lyrics (CER), Whisper performs best compared to the other two. It is assumed that the multilingual ASR tasks and the large-scale pretraining data allow Whisper to model more valid and robust semantic information. Moreover, ContentVec will provide better semantic signals than WeNet, although both are pretrained only on speech. It reveals that the self-supervised pretrained models could be more robust than the classical supervised ASR model. (3) When measuring speaker independence (SIM), the speaker similarity results of the three are comparable and all above 0.9, which is difficult to rank from human perception (see our demo page). This means that when using only semantic-based features for SVC, the three can all be considered speaker-agnostic.
On the other hand, the diverse semantic-based features can be complementary in most cases. For example, from WeNet to Wenet + Whisper, and then to Wenet + Whisper + ContentVec, the results of most metrics are promoted stage by stage. We display a case study of melody modeling in Figure 3. It illustrates that after integrating diverse semantic-based features, the trajectories of the melody between converted audios and ground truth are closer. However, we can also find that using only semantic-based features is hard to model melody adequately (the highest F0CORR in Table 3 is 0.572), appearing the “out of tune” for human hearing. Therefore, introducing explicit melody modeling (such as F0 features) for SVC remains necessary in the present technology context.
4.3 Performance of the DSFF-SVC framework (EQ2)
Base Model Semantic-based Features Recording Studio Setting In-the-Wild Setting
F0CORR ($\uparrow$) F0RMSE ($\downarrow$) CER ($\downarrow$) SIM ($\uparrow$) F0CORR ($\uparrow$) F0RMSE ($\downarrow$) CER ($\downarrow$) SIM ($\uparrow$)
TransformerSVC WeNet 0.849 149.3 15.6% 0.878 0.871 210.0 40.0% 0.865
+ Whisper 0.924 77.2 14.9% 0.881 0.848 183.8 18.7% 0.867
+ Whisper + ContentVec 0.931 75.5 16.2% 0.883 0.857 186.7 23.3% 0.868
VitsSVC WeNet 0.937 175.3 19.1% 0.890 0.919 91.3 57.7% 0.869
+ Whisper 0.945 144.4 17.8% 0.890 0.920 86.9 35.2% 0.869
+ Whisper + ContentVec 0.946 112.9 17.7% 0.886 0.921 79.5 32.3% 0.870
DiffWaveNetSVC WeNet 0.936 55.5 15.8% 0.875 0.901 87.8 60.8% 0.855
+ Whisper 0.943 49.5 15.2% 0.884 0.921 73.6 21.1% 0.865
+ Whisper + ContentVec 0.940 55.2 15.7% 0.884 0.919 79.9 23.3% 0.867
Table 4: Objective Evaluation Results of the proposed DSFF-SVC framework.
To verify the generalization and robustness of the proposed DSFF-SVC framework, we conduct experiments based on both the recording studio and the in-the-wild settings for the three base models. Here we also use the auxiliary prosody features (including F0 and energy), to improve the melody modeling and the expressiveness of the system. The objective and subjective evaluation results can be seen in Table 4 and Table 5. It reveals that: (1) Conducting SVC under in-the-wild setting is more difficult. Facing to the more complex acoustic environment, only using the supervised WeNet produces a very poor performance. Integrating with the weak-supervised Whisper and the self-supervised ContentVec are usually beneficial, especially for Whisper. (2) For the generalization ability, we can observe that for all three models, the proposed idea of diversion semantic-based features fusion is effective in most cases under both settings. (3) For the robustness, particularly, under the in-the-wild setting, the improvement of integrating diverse semantic-based features is more obvious, especially for the objective CER and the subjective metrics.
Compared with Table 3 and 4, there are also many interesting observations about auxiliary prosody features (F0 and energy): (1) After injecting such prosody signals into SVC models, the intelligibility (CER) has been improved a lot. In our opinion, this is because the prosody in singing voice is expressive and characteristic, which is hard to be modeled without the explicit assistant of prosody features like F0. Therefore, using only semantic-based features, some tone- and intonation-related signals will not be learned well, leading to the worse intelligibility. (2) Except for the intelligibility, it is also notable about the change of conversion speaker similarity (SIM). After introducinh F0 features, the melody-related metrics have been improved a lot and the converted singing voices are not out of tune any more. In the meantime, however, the speaker similarity is also harmed. We suppose such auxiliary prosody features own some speaker-specific signals, e.g. the personalized vibrato patterns within F0 features. Therefore, directly copying the original prosody features from the source audio could not be the best choice for SVC, although it is a common way in the most SVC works.
Semantic-based Features Recording Studio In-the-Wild
Nat. ($\uparrow$) Sim. ($\uparrow$) Nat. ($\uparrow$) Sim. ($\uparrow$)
WeNet 2.72 ±$\pm 0.22$ 2.64 ±$\pm 0.21$ 2.85 ±$\pm 0.21$ 2.34 ±$\pm 0.20$
+ Whisper 4.02 ±$\pm 0.18$ 3.13 ±$\pm 0.17$ 3.70 ±$\pm 0.18$ 2.86 ±$\pm 0.23$
+ Whisper + ContentVec 4.14 ±$\pm 0.19$ 3.25 ±$\pm 0.18$ 3.71 ±$\pm 0.18$ 2.82 ±$\pm 0.23$
Table 5: Subjective MOS evaluation results (with 95% confidence interval) of DiffWaveNetSVC. The full scores of Naturalness (Nat.) and Similarity (Sim.) are 5 and 4.
4.4 Performance of the Resolution Transformation based Features Fusion (EQ3)
Fusion Strategy Computational Cost Conversion Quality
RTX ($\uparrow$) RTF ($\downarrow$) F0CORR ($\uparrow$) CER ($\downarrow$) SIM ($\uparrow$)
Cross attention 191.9 2.57 0.896 16.4% 0.871
Resampling 415.0 0.86 0.936 15.8% 0.875
Table 6: The comparison between resampling and cross attention for features fusion.
To verify the performance and efficiency of the resampling for feature fusing, we select the cross attention [21] as the compared method. Specifically, we utilize only WeNet to extract semantic-based features and try different strategies to align its resolution to that of F0 and energy features. We conduct the experiments on DiffWaveNetSVC and follow NaturalSpeech2 [22] to adopt the cross attention for features fusion within the diffusion model.
Except for the quality evaluation for the conversion results, we also measure the computational cost of the both. Assume the duration of the whole training corpus is $d_{train}$, the duration of the evaluation samples to be converted is $d_{infer}$, the time to train an epoch is $t_{train}$, and the time to infer the evaluation samples is $t_{infer}$. We define RTX as $d_{train}/t_{train}$, which can measure the training efficiency. We define RTF as $t_{infer}/d_{infer}$ to measure the inference latency. The computational platform is a single NVIDIA Tesla V100.
The comparison results between resampling and cross attention are displayed in Table 6. We can see that when fusing the features of mismatched time resolutions for the SVC task, resampling is even slightly better than cross attention in terms of conversion quality. Moreover, as a non-learning method, resampling does not introduce any additional parameters and greatly accelerates both training efficiency (RTX) and inference speed (RTF) much.
5 Conclusion and Future Work
This paper investigates three semantic-based pretrained models of supervised, weak-supervised, and self-supervised fashions for SVC. We discover their capabilities including modeling melody and lyrics are diverse and can be complementary. Based on these findings, we propose the SVC framework based on Diverse Semantic-based Features Fusion. The experimental results confirm the effectiveness of our proposed framework, particularly for real-world applications involving in-the-wild data. In addition, this paper raises other questions worth exploring. Is it possible to pretrain one model that can capture both melody and lyrics signals while being speaker-agnostic at the same time? How should we select the pretraining method and construct the pretraining corpus to improve the robustness? We will leave these to our future research.